1. For the following reaction, 14.9 grams of carbon monoxide are allowed to react with 6.94 grams of water .

carbon monoxide(g) + water(l) carbon dioxide(g) + hydrogen(g)

What is the maximum amount of carbon dioxide that can be formed? grams

What is the FORMULA for the limiting reagent?


What amount of the excess reagent remains after the reaction is complete? grams


2. For the following reaction, 9.60 grams of nitrogen monoxide are allowed to react with 9.21 grams of oxygen gas .

nitrogen monoxide(g) + oxygen(g) nitrogen dioxide(g)

What is the maximum mass of nitrogen dioxide that can be formed? grams

What is the FORMULA for the limiting reagent?


What mass of the excess reagent remains after the reaction is complete? grams
1
Expert's answer
2019-10-04T04:40:50-0400

1. CO+H2O= CO2 + H2.

14.9/28: 6.94/18= 0.532142857 : 0.385555556

0.385555556 x 44 = 16.9644445 g of (CO2).

H2O is limiting reagent.

0.146587301 x 28 = 4.10444443 g CO.


2. 2 NO(g) + O2(g) → 2 NO2(g)


9.6/60:9.21/32= 0.16 : 0.2878125.

0.16 x 96= 14.72 g of (NO2).

NO is limiting reagent.

0.1278125 x 32 = 4.09 g of (O2).